OpenStudy (anonymous):

Solve x2 + 4x − 12 = 0.

Nnesha (nnesha):

there are 4+ method to factor quadratic equation for this question AC method (headphone) would easy where you have to find `2` numbers when you `multiply` them you should get product of `AC` and when you `add` or subtract them you should get `middle` term

Nnesha (nnesha):

\[\huge\rm Ax^2+Bx+C=0\] a=leading coefficient b=middle term c= constant term what is a and c in that eqaution ?

Nnesha (nnesha):

equation*

OpenStudy (anonymous):

x and -12 i think

Nnesha (nnesha):

nice leading coefficient is one \[\rm 1x^2+4x-12\] so what two number you should multiply to get -12 but when you add them you should get 4

OpenStudy (anonymous):

am not sure

Nnesha (nnesha):

1 times -12 = -12 right but 1-12 =-11 but we need 4 now what other number you can multiply to get -12 ?

OpenStudy (anonymous):

is it 2 and -6

Nnesha (nnesha):

2 should be negative cus we need POSITIVE 4 -2 times 6 =-12 6-2=4

Nnesha (nnesha):

so factors are (x + number )(x+number)

OpenStudy (anonymous):

so the answer is -2 and -6

Nnesha (nnesha):

no.

Nnesha (nnesha):

not negative 6

Nnesha (nnesha):

wait do you have to solve for x or find factors ?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

so i was right

Nnesha (nnesha):

no. factors are (x+ first number)(x+2nd number) set it equal to zero remember one of them is negativve

Nnesha (nnesha):

(x+6)(x+(-2)) now solve for x

Nnesha (nnesha):

x+(-2)= 0 x= ?